# Thermal Management Design and Key Technology Validation for PandaX Underground Experiment

### v1

@@ 3.1.5 Heating Mode
The PandaX detector vacuum pumping process, the xenon gas recuperation process, and the Photomultiplier Tube (PMT) cryogenic test rewarming process need heat sources. After turning on the heating mode, valves D and F, gear pump B, and electric heater B are activated. Heat is transferred from the ethanol refrigerant to the load until it reaches the set temperature, and the system automatically exits. In heating, the circulating coolant bypass the vessel and evaporator to save energy. The maximum heating temperature is set at  [^3] K, considering the detector’s maximum operating temperature and the ethanol’s boiling point, corresponding to the ethanol vapor pressure of 47 kPa. Ensuring the sealing of the closed circulation tube line can keep the safety. The cooling process after the heating and vacuuming process of the detector, due to the large temperature difference between the detector and the coolant, the indirect output mode is preferred to ensure the the cooling process smoothly.
@@ 3.2 Coolant
It is divided into direct cooling and indirect cooling according to the position relationship of the evaporator and the thermal load during using the cold source. Direct cooling indicates the cold refrigerant exchanges heat with the thermal load in the evaporator directly, whose heat exchange efficiency is higher but lacks of flexibility. Indirect cooling treats the coolant as the intermediate medium, which means the evaporator cools the coolant first, then the cryogenic coolant is pumped to the thermal load for cooling. The so-called coolant was named the second refrigerant, too. Utilizing indirect cooling is convenient for the operation, maintenance, and management of the refrigeration system, as well as the flexible distribution and control of the cooling power. The indirect cooling system can adopt fewer refrigeration units with a larger scale, which is efficient and reliable generally. In the meanwhile, the cooling temperature is stable and reliable because the coolant is also the cold storage medium with large thermal capacity or high thermal inertia. The disadvantage is the refrigeration efficiency reduction compared with the direct cooling for one more heat exchange.
@@
The coolant needs to keep in the liquid phase in the circulation between the evaporator and the thermal load. The anhydrous ethanol is used as the coolant, whose melting point and boiling point are 159 K and 352 K, respectively, and the vapor pressure is 8 kPa at 298 K. The volatility of the ethanol is moderate without stains left after volatilizing so that the pollution of the experimental equipment is avoided. The disadvantage is that ethanol is flammable, with a flash point of 282 K and a spontaneous combustion point of 695 K; so the tightness of the pipeline is critical. Enhancing the ventilation and installing the ethanol sensor to monitor leakage are required to ensure safety.
@@
In general, melting point is measured in slow heating, cooling and static state [17]. In fact, the flow rate of the ethanol coolant in the ACR is large at highly turbulent state with large cooling rate and temperature gradient, so the liquid flow can still be maintained at subcooling conditions, which has two possible reasons: One is the weak strength of the solid ethanol when the temperature is slightly lower than the melting point, which is similar with the negative correlation of the temperature and the strength of the sea ice and metal materials[18],[19]. Another possibility is that the ethanol coolant is in a fluid-ice state, a mixture of liquid ethanol and tiny solid ethanol crystals [20]. When the temperature is close to or slightly lower than the freezing point, the viscosity of ethanol increases obviously, and the rise in the flow resistance results in a change of the flow characteristics. However, the cooling capacity is increased. In the static test, the ethanol at 150 K in the stainless steel containers has a viscosity similar to the honey. The non-freezing property at the overcooling state is conducive to using ethanol as the coolant, and the lowest output temperature of the ACR is about 150 K in the test.
@@ 3.3 Evaporator and Heat Exchanger
The most important technical indicators of a cryogenic refrigerator are the output cooling power and the lowest approachable temperature, which are limited by the performance of the heat exchanger. The increasing cooling power is required to compensate the efficiency reducing of the heat exchanger, and the acquisition cost of the cooling power gets higher as the working temperature decreasing, so the requirement of the heat exchange efficiency is strict [21].
In the previous refrigerator, the coil heat exchanger is arranged as the evaporator in the ethanol coolant vessel, as shown in figure 2A. Refrigerant flows inside the coil while the ethanol coolant is outside. Because of the large vessel section and the small space occupancy of the coil, the flow rate of the coolant is slow during circulation. When the working temperature is much higher than the freezing point of the ethanol, its low viscosity causes the ascendant heat exchange effect. However, the viscosity of the ethanol increases along with the temperature decreasing, and the thickness of the velocity and temperature boundary layers on the coil surface increases. The flow state tends to change from turbulent to laminar, deteriorating the heat exchange effect. Thus, the lowest output temperature of the refrigerator is limited.
The plate heat exchanger is an efficient and compact heat exchanger composed of a series of parallel thin metal plates with corrugated surfaces, which can be considered as the heat exchanger composed of multiple narrow sub-flow channels in parallel with periodic section changes, resulting in high fluid turbulence and effective heat exchange [22]. However, once the coolant is frozen in a sub-channel at the working condition near the freezing point, the flow resistance of the sub-channel increases, and the flow rate decreases, thus accelerating the solidification process further until the sub-channel is blocked and loses the heat exchange ability, and may leave only one effective sub-channel finally. Thus the effective heat exchange area is much less than the design. So the plate heat exchanger is not suitable.
@@
The coaxial heat exchanger is a concentric tube consisting of inner and outer tubes of different diameters coiled into round or oblong shapes, as shown in figure 2B. Compared to the coil heat exchanger, shell and tube heat exchanger, and plate heat exchanger, the casing heat exchanger has only one section of the limited flow channel. The cross-section of the inner tube is a star-shaped coolant channel, as shown in figure 2C. Generally, the ethanol coolant is in a highly turbulent state with high heat exchange efficiency due to the uneven and section-changing tube surface. Because the high-speed turbulence flow strongly impacts the tube wall, the solidified micro-crystals are difficult to deposit and grow, and a large amount of coolant solidification is delayed. Its disadvantage is the relatively large flow resistance [21], and the circulating pump with high-pressure output is required. The output pressure of the gear pump can reach 2 MPa with a maximum suction pressure of 0.09 MPa. The flow resistance of the evaporator increases when the temperature is close to the freezing point of the ethanol. Thus, the gear pump needs to be installed before the evaporator, and the cooling effect of the heat exchanger after the pump fully functions so that the heat exchange reduces the negative impact of the pump heating of the coolant. As a result, the ACR can obtain effective output at lower temperature.
@@ 3.4 Circulation Pump
An impeller centrifugal pump is usually used in this kind of refrigerator for its simple structure and high efficiency, whose principle is to transport energy to the fluid medium using the rotating impeller. However, when the viscosity of the medium increases, the shaft power increases, and the efficiency and the flow decreases. The loss of the mechanical energy transfers into heat energy to heat the medium, and the output flow is affected by the outlet pressure obviously [23]. It was found that increasing the power of the centrifugal pump can not improve the output cooling power of the refrigerator in the previous test because its self-heating effect offsets the cooling ability enhanced by increasing the pump power. The gear pump is a positive displacement pump that pressurizes the medium by squeezing the gears, which can transport the medium with a larger viscosity. Its output flow rate is nearly proportional to the rotary speed, almost without being influenced by the outlet pressure [24]. Because the rated temperature of the refrigerator is around the freezing point of ethanol, and the ethanol viscosity is different from the room temperature, the gear pump is chosen as the circulating pump. Furthermore, a servo motor is used to adjust its speed, which takes a soft start to ensure the reliability of the refrigerator; the rotation acceleration is adjustable with a maximum of 5 $r/s^{2}$. At the same time, the pressure sensor is set at the output of the pump to ensure the pressure is below 0.6 MPa or other setting values, and the pump stops with an alarm when the pressure exceeds 0.8 MPa.
@@ 3.5 Test Results and Analysis
The ACR dimension is 2 m length, 1.25 m width and 1.9 m height. After the development of the ACR, the electrical heater B in figure 1 is used as the virtual thermal load for testing. The ACR peak power consumption is about 50 kw. The lowest output temperature of the ACR is 151 K with an effective output power of 2 kw. However, the stability is not enough for condensing the coolant at the heat exchanger. The stable output temperature is 155 K with an output cooling power of 2.5 kW, The cooling power is 3 kW, 3.6 kW, 4 kW, 4.5 kW, 5 kW, 5.6 kW and 6.2 kW, respectively, corresponding to temperatures of 158.5 K, 163 K, 169 K, 177 K, 195 K, 216 K and 233 K. The performance curve is shown in figure 3, the ACR cooling power is much higher than PTR and GM refrigerator which are listed in table 1.
@@
As shown in figure 3, the refrigeration efficiency of the ACR is near the AL600 refrigerator around 170 K. However, the price per unit of cooling power is significantly lower than AL600. In the applications of PTR and GM refrigerators, the excess cooling power is compensated by electrical heating, and the effective output cooling power is lower than the rated value. The coolant of the ACR has a large heat capacity, and the compressor stops when the coolant is below the setting temperature. The rest of the power of the circulating pump and the control system power is much smaller than the compressor. In the situation where accurate temperature output is required, the precision temperature control method described in section 5 would be used in the ACR to lose a part of the temperature difference, reflecting the decline of refrigeration efficiency. In a word, the refrigeration efficiency of ACR, PTR, and GM refrigerators can not be compared directly; it is more reasonable to be considered comprehensively in the specific system.
@@
The dewar for the PMT cryogenic test is shown in figure 4, with 450 inner diameters and 735mm height. The total weight of the stainless steel interlayer contact with the coolant is about 80kg. The ACR and dewar are connected with a vacuum-insulated hose with a maximum coolant flow rate of 20 L/min. The temperatures of the gas at the top and bottom of the dewar and the PMTs at different locations are measured, respectively. About 150 minutes is taken to cool down from 300 K to 170 K using the direct output cooling mode, while 40 minutes is spent to return to room temperature using the heating mode, as shown in figure 5. Because the heat transfers from the dewar inner wall to PMT depends on air, the large thermal resistance causes an obvious time delay in the cooling and heating process.
@@ 4 Design of Centralized Cooling System (CCS)
As mentioned before, the PandaX experiment requires cooling at multiple locations, CCS is a promising solution. The cold coolant is transported to various thermal payloads by vacuum-insulated pipelines, and the ACR is the operating cold source and liquid nitrogen is back cold source. Similarly, large-scale scientific facilities such as the Beijing Electron-Positron Collider, Large Hadron Collider, China Spallation Neutron Source, and Large Proton Collider have adopted or plan to adopt the CCS with liquid helium temperature zone that provides several kilowatts of cooling power over distances of several kilometers [25].
@@ 4.1 Vacuum-insulated Pipeline design
The length and width of the CJPL-II experimental hall, where the PandaX experiment is located, are about 60 m and 14 m, respectively. To ensure the coverage of the cooling supply, the cooling distance is designed to be 100m, which means the total length of the cooling pipeline is about 200 m. The inner diameter of the pipeline is designed to be 50 mm with vacuum insulation and CF flange connection. The maximum flow rate of the pipeline is about 300 L/min by design. Based on the design of a round-trip pipeline temperature difference of 2 K, the maximum cooling power can reach 20 kW, meeting all the cooling requirements of the far future experimental system. Both the inner pipeline and the insulation layer need to pass the helium mass spectrometry leak detection for safety requirements.
The CCS has about 2000 L of ethanol coolant, which has a large heat capacity and cold storage ability to stabilize the temperature and enhance reliability. The pipeline’s heat leakage is about 1 W/m, and each valve’s heat leakage is about 1W, so the total heat leakage power is less than 300 W, which is acceptable for the CCS with several kilowatts.
In addition, room-temperature ethanol is needed as a heat source for xenon recuperation [13] and the operation of the distillation tower. Due to the temperature difference being at 140 K, the pipe diameter of 20 mm is enough.
@@ 4.2 Multiple Cooling Sources Management
To ensure the reliability of the cooling supply and meet the high-power cooling requirements during stages such as detector pre-cooling and liquid xenon injection, multiple ACRs need to operate simultaneously. A backup ACR that can be put into operation at any time is required, too. Besides the backup ACR, a liquid nitrogen cooling unit is also needed to improve the system’s reliability in case of long-time power failure or other extreme conditions. The multiple cooling sources management strategy is shown in figure 6.
Taking PandaX-4T as an example to illustrate, only one ACR can meet all the cooling power requirements for normal operation, the ACR A in figure 6 is running, valves VA and VB opens, valve VC closes, the coolant is circulating. In the meanwhile, The ACR B and the liquid nitrogen cooling unit are connected to the centralized cooling system as thermal loads with valves VD, VE, VG, and VH open, and the coolant circulates slowly to ensure the low-temperature state and ready to output cooling power at any time. If the thermal load is relatively large, like PanaX-xT, only one ACR is not powerful enough; two or more ACRs should run simultaneously, even combined with the liquid nitrogen cooling unit.
@@
There is about 2000 L coolant in the centralized cooling system, and its heat capacity is about 4×$\times 10^{6}$ J/K, When the ACRs are shut down, the accumulated cooling capacity can be used to maintain the normal operation of the system. The coolant temperature rose about 1K per 800 seconds for a 5 kW thermal load when all the ACRs stopped briefly. This time is sufficient to start the diesel generator as an emergency backup power source or liquid nitrogen cooling unit. The amount of available liquid nitrogen is about 160 L per container of 200 L, which can provide about 6 kWh of cooling power with the 200 J/g vaporization heat. Unlike PTR and GM refrigerators, the UPS with high power and large capacity is not necessary for ACR; a smaller UPS is sufficient for ACR’s controller and pump, which further saves the cost and space of the experimental system and improves reliability.

### v2

@@ 3.1.5 Heating Mode
The PandaX detector requires heat sources for three key processes: vacuum pumping process, xenon gas recuperation, and the Photomultiplier Tube (PMT) cryogenic test rewarming process. Once the heating mode starts, valves D and F, gear pump B, and electric heater B are triggered. Heat is transferred from the ethanol refrigerant to the load until it reaches the set temperature, and the system automatically shuts down. During the heating mode, the circulating coolant bypasses the vessel and evaporator to save energy. The maximum heating temperature is set at 333 K, considering the maximum operating temperature of the detector and the boiling point of the ethanol, corresponding to the ethanol vapor pressure of 47 kPa. Proper sealing of the closed circulation tube line is essential to maintain safety. Following the heating and vacuuming process of the detector, the cooling process is preferred to utilize the indirect output mode, given the large temperature difference between the detector and the coolant, ensuring a smooth the the cooling operation.
@@ 3.2 Coolant
The cooling system can be categorized into direct cooling and indirect cooling based on the positional relationship between the evaporator and the thermal load when implementing the cold source. Direct cooling involves the heat exchange between the cold refrigerant and the thermal load within the evaporator. This method offers higher heat exchange efficiency, however, it lacks flexibility. In contrast, indirect cooling uses a coolant as an intermediate medium. In this system, the evaporator first cools the coolant, which is then circulated to the thermal load for cooling. This coolant is then referred to as the secondary refrigerant. Indirect cooling simplifies the operation, maintenance, and management of the refrigeration system, facilitating a more flexible distribution and control of the cooling power. Moreover, this system often operates with fewer larger refrigeration units, making it more efficient and reliable. Additionally, the cooling temperature is stable and reliable because the coolant is also the cold storage medium with a large thermal capacity or high thermal inertia. However, it is worth noting that indirect cooling may lead to a reduction in refrigeration efficiency compared to direct cooling, as it involves an additional stage of heat exchange.
@@
The coolant must remain in the liquid phase during circulation between the evaporator and the thermal load. Anhydrous ethanol is used as the coolant, with melting point and boiling point of 159 K and 352 K, respectively, and the vapor pressure is 8 kPa at 298 K. The volatility of the ethanol is moderate without stains left after volatilizing hindering the pollution towards the experimental equipment. The risk is that ethanol is flammable, with a flash point of 282 K and a spontaneous combustion point of 695 K. Therefore ensuring the sealing reliability of the ACR is vital. To maintain safety, it is necessary to enhance ventilation and install ethanol sensors to monitor potential leaks.
@@
Generally, melting point measurements are conducted under slow heating, cooling and static state [17]. The flow rate of the ethanol coolant in the ACR is substantial, occurring in a highly turbulent state with a large cooling rate and temperature gradient. Therefore, the liquid flow can be sustained even under sub-cooling conditions, potentially for two reasons: First, the solid ethanol exhibits weak strength when the temperature is slightly lower than the melting point, akin to the negative correlation of the temperature and the strength of the sea ice and metal materials[18],[19]. Secondly, the ethanol coolant is in a fluid-ice state, a mixture of liquid ethanol and tiny solid ethanol crystals [20]. When the temperature is close to or slightly lower than the freezing point, the viscosity of ethanol significantly increases, increasing the flow resistance and changing the flow characteristics. In short, the cooling capacity is increased. In the static test, the ethanol at 150 K contained within stainless steel containers has a viscosity similar to honey. Its non-freezing property in the over-cooling state makes ethanol a suitable coolant, with the lowest output temperature of the ACR recorded at about 150 K during testing.
@@ 3.3 Evaporator and Heat Exchanger
The primary technical indicators of a cryogenic refrigerator include its output cooling power and the lowest attainable temperature, both of which are constrained by the performance of the heat exchanger. An increase in cooling power is essential to compensate for the efficiency reduction of the heat exchanger, and the acquisition cost of the cooling power gets higher as the working temperature decreases, so the requirement of the heat exchange efficiency is strict [21]. In the previously utilized refrigerator, the coil heat exchanger is the evaporator within the ethanol coolant vessel, as illustrated in figure 3A. The refrigerant flows inside the coil while the ethanol coolant is outside. Due to the large cross-sectional area of the vessel and the compact size of the coil, the flow rate of the coolant is relatively slow during circulation. When the working temperature is much higher than the freezing point of the ethanol, its low viscosity causes the ascendant heat exchange effect. However, as the temperature decreases, the viscosity of the ethanol increases, which in turn thickens the velocity and temperature boundary layers on the coil surface increase. This transition causes the flow state to change from turbulent to laminar, deteriorating the heat exchange effect. Thus, the lowest output temperature of the refrigerator is limited.
The plate heat exchanger is a highly efficient and compact heat exchanger comprising a series of parallel thin metal plates with corrugated surfaces. This design is considered as the heat exchanger with a collection of multiple narrow sub-flow channels in parallel with periodic section changes, facilitating high fluid turbulence and effective heat exchange [22]. However, once the coolant begins to freeze in a sub-channel near its freezing point, the flow resistance in that sub-channel increases, leading to a decrease in flow rate. This facilitates the solidification process until the sub-channel is blocked and loses the heat exchange ability, and may leave only one effective sub-channel finally. Consequently, the effective heat exchange area is far less than the original design, making the plate heat exchanger unsuitable for such conditions.
@@
The coaxial heat exchanger is a concentric tube comprising inner and outer tubes of different diameters coiled into round or oblong shapes, as illustrated in figure 1, figure 2 and figure 3B. Compared to the coil heat exchanger, shell and tube heat exchanger, and plate heat exchanger, the casing heat exchanger has only one section of the limited flow channel. The cross-section of the inner tube is a star-shaped coolant channel, as illustrated in figure 3C. Generally, the ethanol coolant is in a highly turbulent state with high heat exchange efficiency due to the uneven and section-changing tube surface. Because the high-speed turbulence flow strongly impacts the tube wall, the solidified micro-crystals are difficult to deposit and grow. However, this design has the drawback of relatively high flow resistance, necessitating the application of a circulating pump with high-pressure output [21]. The gear pump can generate an output pressure of 2 MPa and a maximum suction pressure of 0.09 MPa. The flow resistance of the evaporator increases when the temperature is close to the freezing point of the ethanol. Therefore, the gear pump needs to be installed before the evaporator, allowing the cooling effect of the heat exchanger to fully operate and mitigate the negative impact of pump-induced heating of the coolant. Therefore, the ACR could achieve effective output at lower temperatures.
@@ 3.4 Circulation Pump
An impeller centrifugal pump is typically employed in this kind of refrigerator due to its simple structure and high efficiency. Its operating principle revolves around transferring energy to the fluid medium via a rotating impeller. However, when the viscosity of the medium increases, the shaft power increases, and the efficiency and the flow decrease. The loss of mechanical energy converted into heat, elevating the medium temperature, and the output flow is significantly affected by the outlet pressure [23]. Previous tests revealed that increasing the power of the centrifugal pump could not improve the output cooling power of the refrigerator because its self-heating effect offsets the cooling ability enhanced by increasing the pump power. In contrast, the gear pump is a positive displacement pump, which pressurizes the medium by interlocking the gears, making it capable of handling fluids with higher viscosity. Its output flow rate is nearly proportional to the rotary speed, almost without being influenced by the outlet pressure [24]. Because the rated temperature of the refrigerator is close to the freezing point of ethanol, which has varying viscosity at room temperature, the gear pump is as the preferred choice for the circulating pump. Furthermore, a servo motor is used to adjust the speed of the pump, featuring a soft start to ensure the reliability of the refrigerator. The rotation acceleration is adjustable up to a maximum of 5 $r/s^{2}$. A pressure sensor is installed at the output of the pump to maintain the pressure below 0.6 MPa. Should the pressure exceed 0.8 MPa, the pump will shut down and trigger an alarm.
@@ 3.5 Test Results and Analysis
The ACR dimensions are 2 m length, 1.25 m width and 1.9 m height with a compressor that has an electric power of about 45 kW. Following the development of the ACR, the electrical heater B in figure 2 serves as the virtual thermal load for testing purposes. The ACR peak power consumption is about 50 kw. The lowest output temperature achieved by the ACR is 151 K, corresponding to an effective output power of 2 kw. However, this temperature is not stable enough for condensing the coolant at the heat exchanger. The ACR maintains a stable output temperature of 155 K with an output cooling power of 2.5 kW. Additionally, the cooling power is 3 kW, 3.6 kW, 4 kW, 4.5 kW, 5 kW, 5.6 kW and 6.2 kW, respectively, corresponding to temperatures of 158.5 K, 163 K, 169 K, 177 K, 195 K, 216 K and 233 K. The performance curve shown in figure 4, indicates that the ACR cooling power exceeds that of the PTR and GM refrigerators listed in table 1.
@@
As illustrated in figure 4, the refrigeration efficiency of the ACR is near the AL600 refrigerator around 170 K. However, its price per unit of cooling power is significantly lower than AL600. In the applications of PTR and GM refrigerators, the excess cooling power is compensated by electrical heating, and the effective output cooling power is below the rated value. The coolant of the ACR has a large heat capacity, and the compressor stops when the coolant temperature is below the setting temperature. The rest power of the circulation pump and the control system power is much smaller than the compressor. When accurate temperature output is required, the precision temperature control method described in section 5 would be used in the ACR to lose a part of the temperature difference, reflecting the decline of refrigeration efficiency. In a word, the refrigeration efficiency of ACR, PTR, and GM refrigerators can not be compared directly. It is more reasonable to be considered comprehensively in the specific system.
@@
The dewar for the PMT cryogenic test is illustrated in figure 5, with 450 mm inner diameters and 735 mm height. The total weight of the stainless steel interlayer in contact with the coolant is about 80 kg. The ACR and dewar are connected via a vacuum-insulated hose with a maximum coolant flow rate of 20 L/min. Temperatures are measured at various locations: the gas at the top and bottom of the dewar, and the PMTs. The cooling process, utilizing direct output cooling mode, takes about 150 min to lower the temperature from 300 K to 170 K. In contrast, the return to room temperature via heating mode requires approximately 40 min, as illustrated in figure 6. The heat transfers from the inner wall of the dewar to the PMTs depends on air, leading to large thermal resistance that introduces a noticeable time delay in the cooling and heating process.
@@ 4 Design of Centralized Cooling System (CCS)
As mentioned before, the PandaX experiment requires cooling at multiple locations, CCS is a promising solution. Cold coolant is transported to various thermal payloads via vacuum-insulated pipelines, with the ACR serving as the primary cold source and liquid nitrogen as the back cold source. Similarly, large-scale scientific facilities such as the Beijing Electron-Positron Collider, Large Hadron Collider, China Spallation Neutron Source, and Large Proton Collider have adopted or plan to adopt the CCS with liquid helium temperature zone that provides several kilowatts of cooling power over distances of several kilometers [25]. It should be noted that ethanol coolant is flammable, necessitating using helium mass spectrometry for leak detection in CCS pipelines.
@@ 4.1 Vacuum-insulated Pipeline design
The length and width of the CJPL-II experimental hall, where the PandaX experiment is located, are about 60 m and 14 m, respectively. To ensure cooling coverage distance is designed to reach 100 m, resulting in a total cooling pipeline length of about 200 m. The inner diameter of the pipeline is 50 mm with vacuum insulation and CF flange connection. The maximum flow rate of the pipeline is about 300 L/min by design. Based on the design of a round-trip pipeline temperature difference of 2 K, the maximum cooling power can reach 20 kW, meeting all the cooling requirements of the future experimental system. The inner pipeline and the insulation layer need to pass the helium mass spectrometry leak detection test for safety requirements.
The CCS 2000 L of ethanol coolant has a large heat capacity and cold storage ability to stabilize the temperature and enhance reliability. The heat leakage of the pipeline is about 1 W/m, and the heat leakage of each valve is about 1W. Therefore, the total heat leakage power is below 300 W, a level that is acceptable considering the CCS can handle several kilowatts.
Moreover, room-temperature ethanol is needed as a heat source for xenon recuperation [13] and the operation of the distillation tower. Due to the temperature difference being at 140 K, the pipe diameter of 20 mm is enough.
@@ 4.2 Multiple Cooling Sources Management
To ensure the reliability of the cooling supply and to meet the high-power cooling requirements during stages such as detector pre-cooling and liquid xenon injection, multiple ACRs need to operate simultaneously. Additionally, a backup ACR that can be activated at any time is required. Besides the backup ACR, a liquid nitrogen cooling unit is also required to improve the reliability of the system during prolonged power failures or other extreme conditions. The strategy for managing multiple cooling sources is illustrated in figure 7.
Taking PandaX-4T as an example, a single ACR can meet all the cooling power requirements under normal operation. In this case, ACR A in figure 7 is running, valves VA and VB open, while valve VC closes, allowing the coolant circulate. Meanwhile, ACR B and the liquid nitrogen cooling unit are connected to the CCS as thermal loads with valves VD, VE, VG, and VH open, to enable slow circulation of the coolant, maintaining a low-temperature state and ensuring that cooling power could be delivered at any moment. However, in cases where the thermal load is relatively larger, such as PanaX-xT, a single ACR may not provide sufficient power; hence, it may be necessary to operate two or more ACRs in conjunction with the liquid nitrogen cooling unit.
@@
The CCS contains about 2000 L coolant, with a heat capacity of about 4×$\times 10^{6}$ J/K, When the ACRs are temporarily shut down, the accumulated cooling capacity can be used to maintain the normal operation of the system. During this shutdown, the coolant temperature increases by about 1 K per 800 seconds under a thermal load of 5 kW. This duration is adequate to start the diesel generator as an emergency backup power source or liquid nitrogen cooling unit. The amount of available liquid nitrogen is about 160 L per container of 200 L, which can provide about 6 kWh of cooling power with the 200 J/g vaporization heat. Unlike the PTR and GM refrigerators, a high power and large capacity UPS is not necessary for ACR; a smaller UPS is sufficient to power the controller and pump, leading to cost and space saving for the experimental system, while improving overall reliability.